OpenStudy (anonymous):

Which expressions show how 8 • 54 can be rewritten using the distributive property? Choose all answers that are correct. A. 8 • 50 + 4 B. 8 • 50 + 8 • 4 C. 8 • 60 – 8 • 6 D. 8 • 60 + 8 • 4

OpenStudy (anonymous):

A

OpenStudy (steve816):

Yup it's A

OpenStudy (saurabh135):

B 8*54= 8*(50+4) = 8*50+8*4

OpenStudy (anonymous):

I need more than one answer

OpenStudy (saurabh135):

yup C also

OpenStudy (anonymous):

@Saurabh135 so its c b and a or c and b?

OpenStudy (saurabh135):

its B and C